OpenStudy (kpoprocks21):

Convert the rectangular equation 3x+4y=2 to polar form. I did get an answer, but I need a second opinion please

OpenStudy (kittiwitti1):

http://prntscr.com/c0rvkm

OpenStudy (agent0smith):

First replace x with \(r \cos \theta \) and y with \(r \sin \theta \) then you can factor out r, to solve for it.

OpenStudy (kpoprocks21):

I got 2/ 3cos theta + 4sin theta or is it the other way around?

OpenStudy (agent0smith):

Idk. Show work. \[\large 3 r \cos \theta +4 r \sin \theta=2\]